OpenStudy (anonymous):

Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (–2, – 16), and x intercepts at x = –6 and x = 2 y = x^2 + 4x − ___

OpenStudy (rajee_sam):

(-2,16) means when x is -2 y value is -16. So std form of Quadratic is y = ax^2 + bx + c. We need to find c. put x = -2 and y = -16 and find the value of c. -16 = (-2)^2 + 4(-2) + c; -16 = 4 - 8 + c; -16 = -4 + c; c = -12;